Find the charge q(t) flowing through an element if the current is


i(t) = 3e

−t + 20 cos(15t − π

6

⁄ )uA


q(0) = 1uC


1
Expert's answer
2021-08-18T09:06:01-0400

"i(t)= 3e^{-t}+20\:cos\left(15t-\frac{\pi }{6}\right)\\
q= \int i(t) dt\\
\int \:3e^{-t}+20\cos \left(15t-\frac{\pi }{6}\right)dt\\
\mathrm{Apply\:the\:Sum\:Rule}:\quad \int f\left(x\right)\pm g\left(x\right)dx=\int f\left(x\right)dx\pm \int g\left(x\right)dx\\
=\int \:3e^{-t}dt+\int \:20\cos \left(15t-\frac{\pi }{6}\right)dt\\
=-3e^{-t}+\frac{4}{3}\sin \left(15t-\frac{\pi }{6}\right)\\
=-3e^{-t}+\frac{4}{3}\sin \left(15t-\frac{\pi }{6}\right)+C\\
q(0) = 1uC \implies c=-\frac{4}{3}\\
q=-3e^{-t}+\frac{4}{3}\sin \left(15t-\frac{\pi }{6}\right)-\frac{4}{3}\\"